• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name P. Djawotho

    Log entry time 19:41:59 on October 27,1998

    Entry number 3872

    keyword=target

    17:19 NMR measurement. Target moved to Pick-up coils position: 20023
    17:27 NMR peak is -12.6 mV; EPR/NMR polarization is 35.42/34.11 %
    17:30 Par. lasers off
    17:30-17:36 Rotate field to 270 deg
    17:36 Perp. laser# 4 on
    17:42 Laser# 7 on
    17:45 Laser# 8 on
    17:55 No beam. MCC checks for 45 minutes.
    Keep pumping, but in longitudinal mode.
    Perp lasers off. Moved target to EMPTY: 78184.
    17:58 Rotate back to 0 deg.
    18:03 End of rotation; || laser #1 on.
    18:05 Laser #2 on.
    18:12 Laser #3 on.
    18:21 MCC is ready. Target moved to 3He position: 39246
    18:23 Par. lasers off. Rotate field to 270 deg.
    18:29 Transverse laser #4 on.
    18:31 Laser #7 on.
    18:33 Laser #8 on.